# Deep Linear Probe Generators for Weight Space Learning

### v1

@@ 4.3 Deep Linear Probe Generators
64 Learned Probes 0.873 ±$\pm 0.026$ 0.784 ±$\pm 0.017$ 0.933 ±$\pm 0.001$ 0.885 ±$\pm 0.008$
Synthetic Probes 0.899 ±$\pm 0.022$ 0.832 ±$\pm 0.010$ 0.918 ±$\pm 0.001$ 0.882 ±$\pm 0.007$
128 Learned Probes 0.955 ±$\pm 0.005$ 0.808 ±$\pm 0.006$ 0.936 ±$\pm 0.001$ 0.889 ±$\pm 0.008$
Synthetic Probes 0.959 ±$\pm 0.006$ 0.856 ±$\pm 0.007$ 0.917 ±$\pm 0.001$ 0.893 ±$\pm 0.014$
(a) Latent Optimization
(b) ProbeGen
Figure 3: Learned Probes. Optimized probes by (a) Latent Optimization and (b) ProbeGen for the CIFAR10 Wild Park benchmark. Runs are identical (including seed), except the generator module.
Deep Linear Networks.
Non-linear activations such as ReLU are the engine that makes deep networks very expressive. However, in this case, we would like to add regularization (see Sec. 4.2) rather than increasing expressivity. We therefore remove the activations from our generator, keeping only stacked linear layers. Work by Arora et al. [1] showed that when using SGD, deep linear networks have an implicit regularization effect. We therefore use deep linear networks in our approach, i.e., stacked linear layers but without the non-linear activations between them. In Sec. 5.2, we show that removing the activations reduces overfitting and therefore also enhances performance.
Data-Specfic Inductive Bias.
In the case that our target model $f$ takes structured inputs such as images, we hypothesize that presenting probes from a similar distribution will achieve higher accuracy and reduce overfitting. As we do not know the data distribution, we cannot learn an accurate generative model for it. Still, we can introduce data-specific inductive biases into the generator. E.g., we present a generator for images, where we simply stack 2D convolutional layers, similarly to DCGAN [28]. While this does not guarantee natural image statistics, it encourages some local structure and multi-scale hierarchy, which are some of the most important image characteristics.
5 Experiments
Our evaluation follows the standard protocol for weight space learning. We evaluate on two tasks: training dataset classification and generalization error prediction.
Baselines.
We compare our method, ProbeGen, with StatNN [38] and Neural Graphs [21]. StatNN computes $7$ statistics for the weights and biases of each layer, concatenates them and trains a gradient boosted tree method on this representation. Neural Graphs treat neural networks as computational graphs, where each bias is a node and the weights are the matching edges between these nodes. This method then trains a transformer on the created graph, so the attention score between a pair of neurons depends on the weight that connects them.
Datasets.
We evaluate on $4$ established datasets. For training data prediction we choose the MNIST and FMNIST implicit neural representation (INR) benchmarks [26]. Both datasets were formed by training an INR [37] model for each image of the original dataset. The goal is predicting the class of an image given its INR network. For generalization error prediction, we used the CIFAR10-GS [38] and CIFAR10 Wild Park tasks [21]. These datasets consists of thousands of small CNN models ($3-5$ layers), each trained separately on CIFAR10 [22]. We use accuracy for weight classification and Kendall’s $\tau$ for regression.
Table 3: Results for Small Scale Benchmarks. Comparison of ProbeGen, to graph based, mechanistic approaches and latent optimized probes. We average the results over $5$ different seeds.
Accuracy Kendall’s $\tau$ ($\uparrow$)
# Probes Method MNIST FMNIST CIFAR10-GS CIFAR10 Wild Park
0 StatNN 0.398 ±$\pm 0.001$ 0.418 ±$\pm 0.002$ 0.914 ±$\pm 0.000$ 0.719 ±$\pm 0.010$
Neural Graphs 0.923 ±$\pm 0.003$ 0.727 ±$\pm 0.006$ 0.935 ±$\pm 0.000$ 0.817 ±$\pm 0.007$
64 Neural Graphs 0.967 ±$\pm 0.002$ 0.736 ±$\pm 0.012$ 0.938 ±$\pm 0.001$ 0.888 ±$\pm 0.009$
Vanilla Probing 0.873 ±$\pm 0.026$ 0.784 ±$\pm 0.017$ 0.933 ±$\pm 0.001$ 0.885 ±$\pm 0.008$
ProbeGen 0.980 ±$\pm 0.001$ 0.861 ±$\pm 0.004$ 0.956 ±$\pm 0.000$ 0.933 ±$\pm 0.005$
128 Neural Graphs 0.976 ±$\pm 0.001$ 0.745 ±$\pm 0.008$ 0.938 ±$\pm 0.000$ 0.885 ±$\pm 0.005$
Vanilla Probing 0.955 ±$\pm 0.005$ 0.808 ±$\pm 0.006$ 0.936 ±$\pm 0.001$ 0.889 ±$\pm 0.008$
ProbeGen 0.984 ±$\pm 0.001$ 0.877 ±$\pm 0.003$ 0.957 ±$\pm 0.001$ 0.932 ±$\pm 0.006$
@@ 5.1 Main Results
Accuracy.
Tab. 3 summarizes the results on the standard benchmarks. On INR dataset prediction tasks, ProbeGen outperforms all benchmarks significantly, even when the baselines use the same number of probes. E.g., in the FMNIST class prediction task, ProbeGen improves vanilla probing by at least $6\%$, and outperforms all other approaches by more than $12\%$, reaching an accuracy of $87.7\%$. ProbeGen also outperforms the baselines on predicting the generalization error of CNNs. While vanilla probing performs similarly to graph approaches on CIFAR10-GS, ProbeGen achieves the highest result, and is able to improve the quality of the probes. On CIFAR10 Wild Park ProbeGen also outperforms all previous approaches significantly, improving Kendall’s $\tau$ from $0.889$ to $0.933$.
Table 4: FLOPs Comparison. Tested with $128$ probes and a batch size of $64$.
@@
We compare the computational cost of our method to the baseline in terms of floating point operations (FLOPs). We show the results for MNIST INRs and CIFAR10-GS datasets, having $128$ probes and a batch size of $64$ for both approaches in Tab. 4. We see that probing is a much more efficient approach, requiring between $1.5-3$ orders of magnitude fewer FLOPs.
@@ 5.2 Ablation Studies
By observing Fig. 3, in Sec. 4.3 we hypothesized that removing the activations between the linear layers of the generator will reduce its overfitting. We therefore compare the results with and without non-linear activations. Indeed, as seen in Fig. 5, using non-linear activations results in a higher generalization gap, i.e., more overfitting. The amount of overfitting is even worse when not using a generator at all. In Fig. 5 we see that these activations also harm the model’s performance, suggesting they lead to probes with reduced generalization abilities.
Table 5: Number of Probes. We compare ProbeGen to vanilla probing with different numbers of learned probes. For FMNIST we report accuracy, and for CIFAR10 Wild Park the Kendall’s $\tau$. We average over $5$ different seeds.
FMNIST ($\uparrow$) CIFAR10 Wild Park ($\uparrow$)
# Probes Vanilla ProbeGen Vanilla ProbeGen
16 0.686 ±$\pm 0.043$ 0.805 ±$\pm 0.006$ 0.882 ±$\pm 0.010$ 0.926 ±$\pm 0.007$
32 0.764 ±$\pm 0.013$ 0.844 ±$\pm 0.006$ 0.884 ±$\pm 0.008$ 0.930 ±$\pm 0.005$
64 0.784 ±$\pm 0.017$ 0.861 ±$\pm 0.004$ 0.885 ±$\pm 0.008$ 0.933 ±$\pm 0.005$
128 0.808 ±$\pm 0.006$ 0.877 ±$\pm 0.003$ 0.889 ±$\pm 0.008$ 0.932 ±$\pm 0.006$
Figure 4: Overfitting for Different Generators. We compare ProbeGen to linear generators with fully connected layers, and non-linear convolutional generators. Overfit is measured by the generalization gap of each method. Results are averaged over $5$ seeds.
Figure 5: Ablation Studies. We compare the performance of ProbeGen, linear generators with fully connected layers, and non-linear convolutional generators, for varying numbers of layers. With less than $2$, all versions are equivalent to ProbeGen. We average over $5$ seeds.
@@ 6 Discussion
In this paper we showed the potential of probing methods for learning from models. A side benefit of probing compared to other approaches, is that it is suitable for inference on black-box models without any further adjustments. As probing only observes the responses received from its learned probes, at inference time these could be simply inserted to the black-box model by its API, then making a prediction based on the outputs.
@@
ProbeGen incorporated inductive biases into the probe generator module. We tested it on images, showing its potential when having the right inductive bias for the task. We believe that other modalities could require other structural biases for the probes. E.g., in audio, a consistency term would probably be helpful to simulate realistic recordings, and textual data may even require some pre-training for linguistic priors of the probes. This is left for future work.
@@
One interesting future direction for improving probing is being able to adaptively choose the probes used to test each model. Probing models can than learn a policy for this adaptive selection, which would potentially improve accuracy for a given certain probe budget. An early work by Herrmann et al. [17] showed this could be effective for sequential models.
@@ 7 Limitations
In Sec. 5.1 we showed probing is much more computationally efficient than previous graph based approaches. Still, it requires forwarding the entire model a few times and computing the gradients through the model, with multiple models in each batch. This means that in order to infer about a model using probing, one would need computational resources equivalent to training such a model. This would require a non-trivial solution for learning from larger models, e.g. CLIP [29] or Stable Diffusion [30], under a limited compute budget.
@@ 8 Conclusions
Appendix A Implementation Details
Generator Architectur.
For image generation, we use a transposed convolution based generator. With each layer, the feature maps spatial sizes doubles in each axis (and $4$ times overall), while the number of neurons (channels) decrease by half. We choose a generator width multiplier of $16$, i.e., our generators last layer has $16$ input channels, and the channels multiply by $2$ with each previous layer. Our image generators has $6$ transposed convolutional operators for experiments on the CIFAR10-GS case, and $5$ for experiments on the CIFAR10 WIld Park dataset.
@@ Appendix A Implementation Details
We use a learning rate of $3\cdot 10^{-4}$ and a batch size of $32$ in all our experiments. Our MLP classifier $C$, uses $6$ layers with a hidden size of $256$. The latent vectors of each probe are of size $32$. Except for the CIFAR10-GS [38] dataset, we train all probing algorithms for $30$ epochs. Due to CIFAR10-GS relatively small size, experiments on that dataset were trained for $150$ epochs.
@@
Appendix B Queries of Different Probing Approaches
@@
(a) $2$ Layers with Activations
(b) $3$ Layers with Activations
(c) $4$ Layers with Activations
(d) ProbeGen ($5$ Layers) with Activations
@@
(a) $2$ Fully-Connected Layers
(b) $3$ Fully-Connected Layers
(c) $4$ Fully-Connected Layers
(d) $5$ Fully-Connected Layers

### v2

@@ 5 Experiments
0 StatNN 0.398 $\pm 0.001$ 0.418 $\pm 0.002$ 0.914 $\pm 0.000$ 0.719 $\pm 0.010$
DWS 0.857 $\pm 0.006$ 0.671 $\pm 0.003$ - -
NFNHNP 0.791 $\pm 0.008$ 0.689 $\pm 0.006$ 0.934 $\pm 0.001$ -
ScaleGMNB 0.966 $\pm 0.002$ 0.808 $\pm 0.001$ 0.941 $\pm 0.000$ -
Neural Graphs 0.923 $\pm 0.003$ 0.727 $\pm 0.006$ 0.935 $\pm 0.000$ 0.817 $\pm 0.007$
64 Neural Graphs 0.967 $\pm 0.002$ 0.736 $\pm 0.012$ 0.938 $\pm 0.001$ 0.888 $\pm 0.009$
Vanilla Probing 0.873 $\pm 0.026$ 0.784 $\pm 0.017$ 0.933 $\pm 0.001$ 0.885 $\pm 0.008$
ProbeGen 0.980 $\pm 0.001$ 0.861 $\pm 0.004$ 0.956 $\pm 0.000$ 0.933 $\pm 0.005$
128 Neural Graphs 0.976 $\pm 0.001$ 0.745 $\pm 0.008$ 0.938 $\pm 0.000$ 0.885 $\pm 0.005$
Vanilla Probing 0.955 $\pm 0.005$ 0.808 $\pm 0.006$ 0.936 $\pm 0.001$ 0.889 $\pm 0.008$
ProbeGen 0.984 $\pm 0.001$ 0.877 $\pm 0.003$ 0.957 $\pm 0.001$ 0.932 $\pm 0.006$
@@ 5.1 Main Results
Small-Scale Performance.
Tab. 3 summarizes the results on the standard benchmarks. On INR dataset prediction tasks, ProbeGen outperforms all benchmarks significantly, even when the baselines use the same number of probes. E.g., in the FMNIST class prediction task, ProbeGen outperforms all other approaches by more than $6\%$, reaching an accuracy of $87.7\%$. ProbeGen also outperforms the baselines on predicting the generalization error of CNNs. While vanilla probing performs similarly to graph approaches on CIFAR10-GS, ProbeGen is clearly able to improve the quuality of the probes, and achieves the highest result. On CIFAR10 Wild Park ProbeGen also outperforms all previous approaches significantly, improving Kendall’s $\tau$ from $0.889$ to $0.933$.
Table 4: FLOPs Comparison. Using $128$ probes and a batch size of $64$. ProbeGen is much more efficient than graph approaches.
@@
Table 5: Results for ResNet Scale. ProbeGen can successfully scale to larger sized architectures, and outperforms both baselines.
Kendall’s $\tau$ ($\uparrow$)
# Probes Method ResNet18 Zoo
0 StatNN 0.856
128 Vanilla Probing 0.843
128 ProbeGen 0.910
@@
We compare the computational cost of ProbeGen and Neural-Graphs in terms of floating point operations (FLOPs). We test on the MNIST INRs and CIFAR10-GS datasets, with $128$ probes and a batch size of $64$ for both approaches in Tab. 5. We see that probing is a much more efficient approach, requiring between 1.531.5-3 orders of magnitude fewer FLOPs.
Large models.
Knowing that probing methods are more efficient, we now wish to test if probing methods can really scale to larger size models. We create a new dataset of over $6,000$ ResNet18 He et al. (2016) models. Each ResNet model was trained on a a randomly selected subset of Tiny-Imagenet Le & Yang (2015); Deng et al. (2009). We sampled the subset out of a closed list of $10$ subsets, that we created in advance. For each model, we record its generalization (test) error, and the objective is to predict this error given the model weights. Graph based methods were too computationally expensive in this scale. Tab. 5 shows the results. ProbGen achieves the best results reaching a $0.91$ kendall’s $\tau$, significantly outperforming both vanilla probing and statNN.
@@ 5.2 Ablation Studies
By observing Fig. 3, we hypothesized that removing the activations between the linear layers of the generator will reduce its overfitting. We therefore compare the results with and without non-linear activations. Indeed, as seen in Fig. 5, using non-linear activations results in a higher generalization gap, i.e., more overfitting. The amount of overfitting is even worse when not using a generator at all. In Fig. 5 we see that these activations also harm the model’s performance, suggesting they lead to probes with reduced generalization abilities. In App. A we provide ablations demonstrating that, for these tasks, a deep linear generator outperforms other regularizations.
Figure 4: Overfitting for Different Generators. We compare ProbeGen to linear generators with fully connected layers, and non-linear convolutional generators. Overfit is measured by the generalization gap of each method. Results are averaged over 55 seeds.
Figure 5: Ablation Studies. We compare the performance of ProbeGen, linear generators with fully connected layers, and non-linear convolutional generators, for varying numbers of layers. With less than $2$, all versions are equivalent to ProbeGen. We average over 55 seeds.
@@
Table 6: Number of Probes. We compare ProbeGen to vanilla probing with different numbers of learned probes. We average over 55 different seeds.
FMNIST ($\uparrow$) CIFAR10 Wild Park ($\uparrow$)
# Probes Vanilla ProbeGen Vanilla ProbeGen
16 0.686 $\pm 0.043$ 0.805 $\pm 0.006$ 0.882 $\pm 0.010$ 0.926 $\pm 0.007$
32 0.764 $\pm 0.013$ 0.844 $\pm 0.006$ 0.884 $\pm 0.008$ 0.930 $\pm 0.005$
64 0.784 $\pm 0.017$ 0.861 $\pm 0.004$ 0.885 $\pm 0.008$ 0.933 $\pm 0.005$
128 0.808 $\pm 0.006$ 0.877 $\pm 0.003$ 0.889 $\pm 0.008$ 0.932 $\pm 0.006$
@@ 6 Discussion
We showed the potential of probing for learning from models. A side benefit of probing, compared to other approaches, is that it is suitable for inference on black-box models without any further adjustments. As probing only observes models responses, at inference time it could simply probe a black-box model by its API, then make a prediction based on the outputs.
@@
ProbeGen incorporates inductive biases into the probe generator module. We tested it on images, showing its potential when having the right inductive bias, however, other modalities could require different structural biases. E.g., in audio, a consistency term would probably be helpful to simulate realistic recordings, and textual data may even require some pre-training for linguistic priors of the probes. In App. B we present a preliminary experiment using INRs trained on point clouds, demonstrating that ProbeGen’s success generalizes to other modalities as well.
@@
One interesting future direction for improving probing is being able to adaptively choose the probes used to test each model. Probing models can than learn a policy for this adaptive selection, which would potentially improve accuracy for a given certain probe budget. An early work by Herrmann et al. (2024) showed this could be effective for sequential models.
@@ 7 Limitations
In Sec. 5.1 we showed probing is much more computationally efficient than previous graph based approaches. Still, it requires forwarding the entire model a few times and computing the gradients through the model, with multiple models in each batch. This means that in order to infer about a model using probing, one would need computational resources equivalent to training such a model. This would require a non-trivial solution for learning from larger models, e.g. CLIP (Radford et al., 2021) or Stable Diffusion (Rombach et al., 2022), under a limited compute budget.
@@ 8 Conclusions
9 Social Impact
The goal of this paper is to further our understanding of neural networks. We expect this increased understanding to be helpful for reducing the social risks of AI such as bias and model safety.
10 Acknowledgments
This work was partially supported by the Israel Science Foundation (ISF), the Council for Higher Education (Vatat), the Center for Interdisciplinary Data Science Research (CIDR), the Israeli Cyber Authority, and KLA.
Appendix A Other Regularizations
We compare deep linear generators to a range of $\ell_{2}$ weight decay regularization strengths. We present the results in Tab. 7. Our results show that while adding some regularization might be helpful, it is not as good as a deep linear generator, and requires aggressive hyper-parameter tuning of the regularization strength, including non-standard values (a weight decay of $10^{-7},10^{-8}$), where a deep linear generator does not. Moreover, these regularizations are not consistent between experiments, and there is no single value that fits all experiments. We conclude that a deep linear generator, as presented in ProbeGen, provides a stable and subtle regularization that fits weight-space analysis.
Table 7: Performance Comparison of Different Regularizations vs. ProbeGen.
Accuracy Kendall’s $\tau$ ($\uparrow$)
Method MNIST FMNIST CIFAR10-GS CIFAR10 Wild Park
Non-Linear Gen. + L2 Reg. ($10^{-2}$) 0.942 0.531 0.758 0.442
Non-Linear Gen. + L2 Reg. ($10^{-3}$) 0.979 0.858 0.824 0.547
Non-Linear Gen. + L2 Reg. ($10^{-4}$) 0.981 0.867 0.861 0.737
Non-Linear Gen. + L2 Reg. ($10^{-5}$) 0.982 0.869 0.911 0.871
Non-Linear Gen. + L2 Reg. ($10^{-6}$) 0.982 0.870 0.941 0.905
Non-Linear Gen. + L2 Reg. ($10^{-7}$) 0.982 0.869 0.947 0.922
Non-Linear Gen. + L2 Reg. ($10^{-8}$) 0.981 0.870 0.948 0.928
ProbeGen 0.984 0.877 0.957 0.932
Figure 8: Ablating the Number of Probes used for Point Cloud INRs Classification. We show that ProbeGen exhibits a much better scaling law when classifying point clouds, compared to Vanilla Probing. Notably, ProbeGen with $32$ probes already outperforms Vanilla Probing with $2048$ probes.
Appendix B Another Data Modality: Point Clouds
We extend our evaluation to include a more complex data modality: point clouds. Specifically, we use the Neural-Field-Arena (Papa et al., 2024), to evaluate ProbeGen’s ability in classifying INRs trained on point clouds from the ShapeNet (Chang et al., 2015) dataset. The goal is to detect the class of the point cloud each INR was trained on, directly from the INR’s weights. As shown in Table 8, ProbeGen consistently outperforms all prior approaches, including Vanilla Probing.
Figure 8 further examines the efficiency of probing methods with respect to the number of probes used. We can see that even under a highly restrictive budget of $32$ probes, ProbeGen is able to surpass all baselines. Notably, this $32$ probes version of ProbeGen outperforms Vanilla Probing with $2048$ probes, showing ProbeGen’s better scaling laws.
Table 8: Point Cloud INRs Classification. We show that ProbeGen can successfully learn from models trained on more complex data modelities, such as point clouds. Number of probes is in parentheses.
Accuracy
Method Point Clouds INRs
DWS 0.911
Neural Graphs 0.903
Vanilla Probing (128) 0.913
ProbeGen 0.930
(a) Vanilla Probing (Latent Optimization)
(b) Single Layer Generator
(c) 33 Layers Generator
(d) ProbeGen (55 Layers)
Figure 9: Learned Queries by Deep Linear Generators. Queries learned by deep linear generators and vanilla probing, for the CIFAR10 Wild Park dataset.
Appendix C Implementation Details
Generator Architecture.
For image generation, we use a transposed convolution based generator. With each layer, the feature maps spatial sizes doubles in each axis (and $4$ times overall), while the number of neurons (channels) decrease by half. We choose a generator width multiplier of $16$, i.e., our generators last layer has $16$ input channels, and the channels multiply by $2$ with each previous layer. Our image generators has $6$ transposed convolutional operators for experiments on the CIFAR10-GS and ResNet18 cases, and $5$ for experiments on the CIFAR10 Wild Park dataset.
@@ Appendix C Implementation Details
We use a learning rate of $3\cdot 10^{-4}$ and a batch size of $32$ in all our experiments. Our MLP classifier $C$, uses $6$ layers with a hidden size of $256$. The latent vectors of each probe are of size $32$. We trained all probing algorithms on the INR and CIFAR10 Wild Park experiments for $30$ epochs, all experiments on the CIFAR10-GS dataset for $150$ epochs, and all experiments on our ResNet18 Model Zoo dataset for $100$ epochs. Additionally, we trained all baselines using their default hyper-parameters.
@@
Appendix D Queries of Different Probing Approaches
@@
(a) 22 Layers with Activations
(b) 33 Layers with Activations
(c) 44 Layers with Activations
(d) ProbeGen (55 Layers) with Activations
@@
(a) 22 Fully-Connected Layers
(b) 33 Fully-Connected Layers
(c) 44 Fully-Connected Layers
(d) 55 Fully-Connected Layers